OpenStudy (anonymous):

Use the elimination method to solve the system. 0.18f-0.3g=3 0.15g-0.9f=-5.55

OpenStudy (anonymous):

@pengembara_bumi1

OpenStudy (uri):

Arrange the second equation first. and multiply first eq with 0.5

OpenStudy (anonymous):

arrange in the same order as the first one

OpenStudy (uri):

0.15g-0.9f=-5.55 0.9f+0.15g=5.55 <---like this so now we have this 0.18f-0.3g=3 (multiply by 0.5) 0.9f+0.15g=5.55

OpenStudy (anonymous):

f = –7, g = 5 f = –5, g = –7 f = 5, g = 7 f = 5, g = –7

OpenStudy (anonymous):

f=-5 ,g=-7